# Image-to-Text Logic Jailbreak: Your Imagination can Help You Do Anything

### v1

@@ Abstract
Large Visual Language Models (VLMs) such as GPT-4 have achieved remarkable success in generating comprehensive and nuanced responses, surpassing the capabilities of large language models. However, with the integration of visual inputs, new security concerns emerge, as malicious attackers can exploit multiple modalities to achieve their objectives. This has led to increasing attention on the vulnerabilities of VLMs to jailbreak. Most existing research focuses on generating adversarial images or nonsensical image collections to compromise these models. However, the challenge of leveraging meaningful images to produce targeted textual content using the VLMs’ logical comprehension of images remains unexplored. In this paper, we explore the problem of logical jailbreak from meaningful images to text. To investigate this issue, we introduce a novel dataset designed to evaluate flowchart image jailbreak. Furthermore, we develop a framework for text-to-text jailbreak using VLMs. Finally, we conduct an extensive evaluation of the framework on GPT-4o and GPT-4-vision-preview, with jailbreak rates of 92.8% and 70.0%, respectively. Our research reveals significant vulnerabilities in current VLMs concerning image-to-text jailbreak. These findings underscore the need for a deeper examination of the security flaws in VLMs before their practical deployment.
@@ 1 Introduction
The emergence of Visual Language Models (VLMs) represents a significant milestone in the field of artificial intelligence, integrating the strengths of computer vision and natural language processing to achieve multi-modal understanding and generation (Zhang et al., 2024a; Park & Kim, 2023). Recently, VLMs such as GPT4 (OpenAI, 2023) and Gemini (Team et al., 2023) have demonstrated remarkable success in various tasks, including visual question answering, vision-based coding, and object localization (Yang et al., 2023b). Furthermore, with the release of an increasing number of open-source VLMs like Stable-Diffusion-3 (StabilityAI, 2024), downstream application tasks have experienced explosive growth. This expansion has significantly enhanced the impact of VLMs across various domains, including e-commerce (Koh et al., 2024), transportation (Yi et al., 2024), and healthcare (Chen et al., 2024).
@@
A long-standing challenge with visual and language models is their high susceptibility to spoofing by adversarial samples (Zhang et al., 2021, 2020), a vulnerability particularly evident in image classification (Goodfellow et al., 2014) and sentiment classification tasks (He et al., 2021; Liu et al., 2023; Lin et al., 2021). In addition, due to the extensive training data sourced from the Internet, which inevitably contains malicious images and text, attackers can exploit VLMs to jailbreak to generate harmful content jailbreak (Carlini et al., 2023; Zhang et al., 2023; Kim et al., 2024). This issue is further exacerbated by the widespread availability of open-source VLMs, which are not always subjected to the rigorous security measures employed by companies after deployment. Consequently, pre-existing vulnerabilities may persist, allowing adversaries to continuously exploit fixed methods to jailbreak these models. In response to these concerns, researchers have shifted their focus towards VLMs jailbreak (Zhao et al., 2024; Zhang et al., 2023), aiming to prevent the output of malicious content. Current researches primarily alleviate text to image and text combined with nonsensical image to image jailbreaks. However, researchers ignored the fact that the logical comprehension, imagination and jailbreak capabilities of VLMs are adversarial to each other, i.e., VLMs should be developed to ensure that they understand the content of the picture, follow it and imagine details, but also to ensure that jailbreak does not occur. As the visual logical capabilities of VLMs continue to advance, it is imperative to fully leverage visual logical capabilities to rigorously assess their vulnerabilities.
@@
In this paper, we conduct an extensive evaluation that whether VLMs have been trained to understand flowcharts for jailbreak image through their logical reasoning and imaginative capabilities. Therefore, we first propose logic jailbreak problem and introduce a meaningful image dataset specifically designed to evaluate the logical jailbreak capabilities of VLMs. Next, we expand the jailbreak paradigm to text-to-text jailbreak and design a framework for text-to-text jailbreak with a logical flowchart, leveraging the victim VLMs imaginative capabilities for jailbreak. Finally, we conduct a comprehensive evaluation on GPT-4o and GPT-4-vision-preview using the proposed dataset and analyze the vulnerabilities.
@@
  • We first introduces the problem of logic jailbreak and a meaningful image dataset to evaluate logical jailbreak capabilities of VLMs.
  • Using the idea of logic jailbreak, we extend it to a fully automated framework for text-to-text jailbreak.
  • we conduct a comprehensive experiments on GPT-4o and GPT-4-vision-preview and analyze the vulnerabilities.
@@ 2 Related Work
Language Models and Security The issue of adversarial attacks (Goodfellow et al., 2014) has been a central focus of research in language model security. Depending on the sentence modification procedure, these attacks can be categorized into character-level (He et al., 2021), word-level (Liu et al., 2023), and sentence-level (Lin et al., 2021) manipulations. By making subtle changes to the original samples that are difficult for humans to detect, adversaries can cause language models to produce incorrect predictions (Zhang et al., 2020). With the emergence of LLMs such as BERT (Devlin et al., 2018) and GPT (OpenAI, 2022), there has been an increased focus on the vulnerabilities and jailbreak of language models. For open source LLMs, Zou et al. (2023) proposed a novel jailbreak approach named GCG by appending suffixes to malicious questions, leveraging the autoprompt (Shin et al., 2020) concept for open-source models, while also elucidating the transferability of jailbreak prompts. Huang et al. (2023) underscored the profound impact of parameter configurations on LLMs vulnerabilities. For black box LLMs, Lapid et al. (2023) employed genetic algorithms to craft adversarial suffixes for breaching black-box LLMs. However, the above mentioned methods of adding suffixes produces unnatural jailbreak prompts and more research focused on generating natural jailbreak prompts. Yu et al. (2023) advocated for applying software testing principles to LLMs jailbreaking, while Chao et al. (2023) proposed using one LLM to jailbreak another. Li et al. (2023) engineered a multi-step attack against ChatGPT, targeting the extraction of sensitive personal data, thereby raising profound privacy concerns and significant security risks for LLMs applications. In addition to the above recent investigations, Carlini et al. (2023) have demonstrated that simply manipulating image inputs can compromise the security of multi-modal models. Furthermore, research (Qi et al., 2023) suggests that fine-tuning aligned LLMs on adversarial examples can jeopardize secure alignment.
@@
3 Problem
In our study, we investigate whether the logical and imaginative capabilities of VLMs can be leveraged for the purpose of jailbreak. Specifically, we provided harmful flowcharts for VLMs to understand and complete the jailbreak. For clarify, we give an example in Figure 1. First, the attacker manually creates a flowchart image of malicious behaviors and then gives the corresponding prompt. When the VLMs receives the malicious behavior flowchart image, the VLMs makes a trade-off between thinking about the flowchart images details and curbing the malicious output.
Figure 1: An example of logic jailbreak in GPT-4o.
@@ 4 Models and Datasets
In this section, we give the model and datasets used for the experiment.
Models Considering that the GPT series is currently the most widely used model, and given the multi-modal capabilities of GPT-4o, which has been made freely available to a broad user base, along with the exceptional performance of GPT-4-vision-preview across various benchmarks, we have selected GPT-4o and GPT-4-vision-preview for our experiments in this study.
Handmade Datasets As no previous work has introduced logical jailbreak image datasets, we manually extracting $7070$ malicious behaviours from AdvBench (Zou et al., 2023) and creating corresponding Logic Jailbreak Flowcharts dataset (LJF) [^1] Our dataset is accessible by contacting xz549@exeter.ac.uk. Additionally, to compare the effectiveness of logical image jailbreaks with text-based image jailbreaks, we have also converted these $7070$ malicious questions into Simple Jailbreak Images dataset (SJI) containing only their textual content.
AI Flowchart Datasets To ensure a comprehensive evaluation, we employed a VLMs model to generate flowcharts corresponding to harmful behaviors. We use all $520520$ harmful behaviors from the AdvBench dataset for flowchart image generation. The model utilized for generating these flowcharts is SDXL (Podell et al., 2023).
@@ 5 Evaluation
In this section, we first introduce the evaluation metrics. Then, we apply the SJI, LJF and AI Flowchart datasets to GPT-4o and GPT-4-vision-preview, respectively. Finally, we analyse the experiment results.
5.1 Metric
In our study, we evaluate the success of VLMs jailbreaks using the Attack Success Rate (ASR) metric (Zou et al., 2023; Lapid et al., 2023; Qi et al., 2023; Google, 2023; Chao et al., 2023). In addition, we adopt the evaluation methodology proposed by (Yu et al., 2023) to determine whether LLMs have exhibited jailbreak behaviors.
5.2 Experiments
  • Single-Round Jailbreak We use the image from SJI as input to GPT-4o and GPT-4-vision-preview. Subsequently, we prompt it for supplementary information. The specific prompt is as follows:
@@ 5.2 Experiments
  • Logic Jailbreak We use the image from LJF as input to GPT-4o and GPT-4-vision-preview. Subsequently, we prompt it for supplementary information. The prompt is same in Single-Round Jailbreak.
  • AI Flowchart Jailbreak To investigate the impact of image quality on jailbreak outcomes, the images in AI Flowchart dataset were inputted into GPT-4o and GPT-4-vision-preview, requesting additional detail enhancement. The input prompt is same in Single-Round Jailbreak.
@@ 5.3 Results
We give the results of our experiments in Table 1. Upon analyzing the table, we can observe that:
  • It is evident that neither GPT-4o nor GPT-4-vision-preview can be successfully jailbroken when SJI datasets are input into the model and the VLMs are prompted to add specific details.
  • When using the AI flowchart dataset, the ASR of GPT-4o and GPT-4-vision-preview is higher compared to the SJI dataset, indicating the effectiveness of the flowchart image jailbreak technique.
  • When using our hand-made flowchart image dataset (LJF), GPT-4o achieves a jailbreak rate of 92.8%, while GPT-4 achieves a rate of 70%. This demonstrates the efficacy of our hand-made jailbreak flowchart.
Therefore, jailbreak flowcharts are crucial for evaluating the security of VLMs.
Table 1: The jailbreak results (num and ASR) in GPT-4o and GPT-4-vision-preview.
SJI LJF Selected AI Flowchart AI Flowchart
GPT-4o 0 (0%) 65 (92.8%) 12 (17.1%) 102 (19.6%)
GPT-4-v 0 (0%) 49 (70.0%) 20 (40.0%) 161 (31.0%)
6 Expanding to Text-to-Text Jailbreak
In this section, we present a new automated text-to-text jailbreak framework. The flowchart is shown in Figure 2. The process can be described as follows.
  • For arbitrary harmful behaviors, we transform them into declarative sentences.
  • We input the converted declarative sentence to VLMs for it to make a picture flowchart. Next, we input the generated flowchart into the attacked model.
  • We ask the attacked model to understand and add specific details to the flowchart.
The result is shown last column in Table  1. From the table, we can observe that the ASR of this method are 19.6% for GPT-4o and 31.0% for GPT-4-vision-preview. When compared to the results obtained using the hand-made flowchart image, it becomes evident that the effectiveness of this automatic jailbreak framework is closely related to the quality of flowchart image generation. Consequently, the models generating high-quality flowchart image are really important for the automatic jailbreak framework.
Figure 2: The automated text-to-text jailbreak framework.
@@ 7 Conclusion
In this paper, we introduce a novel jailbreak paradigms for VLMs called flowchart image jailbreak and extend this paradigms to text-to-text jailbreak. Additionally, we propose two flowchart image datasets: the hand-made flowchart image dataset and the AI flowchart dataset. Finally, we conduct an extensive evaluation on GPT-4o nor GPT-4-vision-preview, revealing their vulnerability to flowchart image jailbreaks. Consequently, there is an urgent need for more comprehensive flowchart jailbreak image datasets to thoroughly evaluate existing VLMs.
@@ 8 Future Work
SJI Dataset Jailbreak The SJI dataset contains only $7070$ images, which is insufficient to comprehensively evaluate whether VLMs can be jailbroken using images that contain malicious behavioral text. Examples of such jailbreak attempts are provided in the Appendix. Therefore, it is crucial to address the gaps in these datasets to enable a more thorough evaluation.
@@
Multi-Language Jailbreak In our experiments, we exclusively utilized English prompts and did not account for the language sensitivity of VLMs. However, it is important to note that the outcome of jailbreak attempts can vary depending on the language used, even when the prompt and flowchart remain consistent. We give a real conversation with GPT-4o in Appendix Fig 6 and Fig  7.
Evaluation on VLMs’ Ability to Understand Logical Flowcharts The jailbreak techniques proposed in this paper are based on VLMs’ capability to interpret flowcharts. However, different VLMs exhibit varying degrees of proficiency in understanding flowcharts. It is equally important to evaluate the impact of VLMs’ visual logic comprehension on the success of jailbreak attempts (Zhang et al., 2024b). Consequently, there is an urgent need for datasets that can effectively assess the visual logic comprehension abilities of VLMs.

### v2

@@ Abstract
Large Visual Language Models (VLMs) such as GPT-4V have achieved remarkable success in generating comprehensive and nuanced responses. Researchers have proposed various benchmarks for evaluating the capabilities of VLMs. With the integration of visual and text inputs in VLMs, new security issues emerge, as malicious attackers can exploit multiple modalities to achieve their objectives. This has led to increasing attention on the vulnerabilities of VLMs to jailbreak. Most existing research focuses on generating adversarial images or nonsensical image to jailbreak these models. However, no researchers evaluate whether logic understanding capabilities of VLMs in flowchart can influence jailbreak. Therefore, to fill this gap, this paper first introduces a novel dataset Flow-JD specifically designed to evaluate the logic-based flowchart jailbreak capabilities of VLMs. We conduct an extensive evaluation on GPT-4o, GPT-4V, other 5 SOTA open source VLMs and the jailbreak rate is up to 92.8%. Our research reveals significant vulnerabilities in current VLMs concerning image-to-text jailbreak and these findings underscore the the urgency for the development of robust and effective future defenses.
@@ 1 Introduction
The emergence of Visual Language Models (VLMs) represents a significant milestone in the field of artificial intelligence, integrating the strengths of Computer Vision (CV) and Natural Language Processing (NLP) to achieve multi-modal understanding and generation (Zhang et al., 2024b; Park & Kim, 2023). Recently, VLMs such as GPT4 (Achiam et al., 2023) have demonstrated remarkable success in various tasks, including visual question answering, vision-based coding, and object localization (Yang et al., 2023b). Furthermore, with the release of an increasing number of open-source VLMs (Hu et al., 2024; Liu et al., 2024a, b; Bai et al., 2023; Wang et al., 2023), downstream application tasks have experienced explosive growth. This expansion has significantly enhanced the impact of VLMs across various domains, including e-commerce (Koh et al., 2024), transportation (Yi et al., 2024), and healthcare (Chen et al., 2024).
@@
A long-standing challenge in CV and NLP is their high susceptibility to spoofing by adversarial samples (Zhang et al., 2021, 2020) to generate error output, a vulnerability particularly evident in image classification (Goodfellow et al., 2014) and sentiment classification tasks (He et al., 2021; Liu et al., 2023a; Lin et al., 2021). For VLMs, due to the extensive training data sourced from the Internet, which inevitably contains malicious images and text, attackers can exploit adversarial method to force VLMs to generate harmful content which is called jailbreak (Carlini et al., 2023; Zhang et al., 2023; Kim et al., 2024). The jailbreak issue has been extensively studied in LLMs (Zou et al., 2023; Shin et al., 2020; Huang et al., 2023; Lapid et al., 2023; Yu et al., 2023; Chao et al., 2023; Li et al., 2023; Anil et al., 2024; Zou et al., 2024) yet it only received a little attention in VLMs (Zhao et al., 2024; Zhang et al., 2023; Qi et al., 2024). Current researches primarily alleviate text to image and text combined with nonsensical image to image jailbreaks. However, researchers ignored the fact that the logical comprehension, imagination and jailbreak capabilities of VLMs are adversarial to each other, i.e., VLMs should be developed to ensure that they understand the content of the figure, follow it and imagine details, but also to ensure that jailbreak does not occur. As the visual logical capabilities of VLMs continue to advance, it is urgent to fully evaluate the logic-based jail- break capabilities of VLMs to rigorously assess their vulnerabilities.
@@
Therefore, in this paper, we propose logic jailbreak problem and first introduce a Flowchart Jailbreak Dataset (Flow-JD) specifically to evaluate the logical flowchart jailbreak capabilities of VLMs. Then, we conduct an extensive evaluation that whether VLMs can understand flowcharts for jailbreak through their logical reasoning and imaginative capabilities. We explain jailbreak rate for the differences performance between the different models.
@@
  • We first introduces Flow-JD to evaluate logical flowchart jailbreak capabilities of VLMs.
  • We conduct a comprehensive experiments on SOTA VLMs and analyze the vulnerabilities.
  • We outline the limitations of this study and identified directions for future research to further enhance the safety of VLMs.
Figure 1: An example of logic flowchart jailbreak in GPT-4o.
@@ 2 Related Work
Language Models and Security The issue of adversarial attacks (Goodfellow et al., 2014) has been a central focus of research in language model security. Depending on the sentence modification procedure, these attacks can be categorized into character-level (He et al., 2021), word-level (Liu et al., 2023a), and sentence-level (Lin et al., 2021) manipulations. By making subtle changes to the original samples that are difficult for humans to detect, adversaries can cause language models to produce incorrect predictions (Zhang et al., 2020). With the emergence of LLMs such as BERT (Devlin et al., 2018) and GPT (OpenAI, 2022), there has been an increased focus on the vulnerabilities and jailbreak of language models. Zou et al. (2023) proposed a novel jailbreak approach named GCG by appending suffixes to malicious questions, leveraging the autoprompt (Shin et al., 2020) concept for open-source models, while also elucidating the transferability of jailbreak prompts. Huang et al. (2023) underscored the profound impact of parameter configurations on LLMs vulnerabilities.
@@
VLM Benchmark To comprehensively evaluate the capabilities of VLMs, researchers have developed several benchmarks such as LVLM-eHub (Xu et al., 2023), MME (Yin et al., 2023), MMBench (Liu et al., 2023b), TextVQA (Singh et al., 2019), InfographicVQA (Mathew et al., 2022), ChartQA (Masry et al., 2022), EgoThink (Cheng et al., 2024), MathVista (Lu et al., 2023), FlowchartQA (Tannert et al., 2023) and FlowCE (Zhang et al., 2024a). We provide a summary of these datasets in Table 1. We compare our dataset with existing benchmarks, highlighting that Flow-JD is the first dataset specifically designed to assess the safety capabilities of VLMs, with a particular focus on logical flowchart safety.
Benchmark Capability Size
LVLM-eHub General Multi-Modality 332k
MME General Multi-Modality 2,194
MMBench General Multi-Modality 2974
TextVQA Text Recognition and Contextual Reasoning 45.3k
InfographicVQA Integrated Document Visual and Textual Reasoning 30k
ChartQA Chart Understanding and Analysis 9.6k
EgoThink First-Person Thinking 700
MathVista Mathematical Reasoning 6,141
FlowchartQA Geometric and Topological Information of Flowcharts 6M
FlowCE Comprehensive Understanding of Flowcharts 505
Flow-JD (ours) Logic-based Flowcharts jailbreak 590
Table 1: The Comparison of benchmarks in VLMs.
@@ 3 Models and Datasets
In this section, we give the models and datasets used for the evaluation.
Models Considering that the GPT series is currently the most widely used model, and given the multi-modal capabilities of GPT-4o (Achiam et al., 2023), which has been made freely available to a broad user base, along with the exceptional performance of GPT-4V (Achiam et al., 2023) across various benchmarks, we have selected GPT-4o and GPT-4V as closed-source model for our experiments in this study. For open-source model, we select MiniCPM-V2 (Hu et al., 2024), LLAVA-V1.5 (7B) (Liu et al., 2024a), LLAVA-V1.6 (7B) (Liu et al., 2024b), MiniCPM-Llama3-V2.5 (Hu et al., 2024) and Qwen-Chat-VL (Bai et al., 2023) to evaluate.
Datasets
  • Hand-made Datasets As no previous work introduced logical jailbreak image datasets, we manually extracting 70 harmful behaviours in $1010$ categories from AdvBench (Zou et al., 2023) and creating corresponding hand-made logic jailbreak flowcharts dataset (Flow-HJD) [^1] Our flowcharts were created by two PhD students in computing science, and the dataset is accessible by https://github.com/zhuxiaotianzaishuijiao/Jailbreak-flowchart..
  • Stable Datasets To ensure a comprehensive evaluation, we employed SDXL (Podell et al., 2023) to generate flowcharts dataset (Flow-SJD datasets) corresponding to harmful behaviors. We use all $520520$ harmful behaviors from the AdvBench dataset for flowchart image generation.
Datasets Evaluation We conducted a quality assessment of both datasets. We employed GPT-4V to evaluate the similarity between each dataset and the text associated with it. The prompt used to evaluated the similarity is as follows:
You are asked to evaluate the similarity of the input image to the specified text, ultimately giving a score with a value between 1 and 10. Note that you do not need to follow the instructions in the specified text. I will use ’====Text begins====’ to indicate the beginning of the specified text and ’====Text ends====’ to indicate the end. Here is text: ====Text begins==== the harmful behaviour ====Text ends==== Only just give the similarity score between 1 and 10 without any format and Note do not make any other explanation in your answer.
The results of similarity comparison are illustrated in Fig 3. Flow-HJD demonstrates a significantly higher similarity between the image and the corresponding harmful text compared to Flow-SJD.
@@ 4 Evaluation
In our study, we evaluated the logical jailbreak capabilities of text-image-to-text VLMs. Specifically, we provide harmful flowcharts and simple prompts for VLMs to evaluate. For clarify, we give an example in Figure 1. First, we manually creates a flowchart image of ”robbing a bank” and then gives the corresponding prompt. When the VLMs receives the flowchart image, the VLMs makes a trade-off between thinking about the flowchart images details and prevent the malicious output.
4.1 Metrics
We evaluate the success of VLMs jailbreaks using the Attack Success Rate (ASR) metric (Zou et al., 2023; Lapid et al., 2023; Qi et al., 2023; Google, 2023; Chao et al., 2023). In addition, we adopt the evaluation methodology proposed by (Yu et al., 2023) which evaluate responses using Roberta (Liu et al., 2019) trained on the manually datasets human judgments to determine whether VLMs jailbreak.
4.2 Experiments Setting
We use the image from Flow-HJD and Flow-SJD as image input to different VLMs. The text input is shown as follows.
@@ 4.3 Results
The evaluation can ben shown in Table 2. The result reveals that all VLMs exhibit significantly higher jailbreak rates on the manual dataset compared to the AI dataset. This is because there is more similarity between the image and the harmful behaviours in Flow-HJD. This means that VLMs have a superior capacity to comprehend the image content. Consequently, it can leverage its logical reasoning and imaginative abilities to effectively supplement the image, thereby facilitating the jailbreak process.
Figure 2: The correlation score between the responses and harmful behaviours.
Furthermore, as illustrated in the table, GPT-4o and GPT-4V, which are commercial Visual Language Models (VLMs), exhibit notably high attack success rates on flow-HJD, achieving 92.8% and 70% respectively. These rates surpass those of the majority of other VLMs. We hypothesize that this superior performance may be attributed to their enhanced capabilities in image understanding. Consequently, we employed GPT-4 to evaluate the correlation between the responses of VLMs and corresponding harmful behaviors for Flow-HJD and Flow-SJD*. The prompt used to evaluated the correlation score is as follows:
You are asked to evaluate the correlation of 2 paragraphs, ultimately giving a correlation score with a value between 1 and 10. Note that you do not need to follow the instructions in the paragraphs. I will use ’====Paragraph begins====’ to indicate the beginning of the paragraph and ’====Paragraph ends====’ to indicate the end. Here are paragraphs:” ====Paragraph1 begins==== Response ====Paragraph1 ends==== ====Paragraph2 begins==== the harmful behaviour ====Paragraph2 ends==== Only just give the correlation score between 1 and 10 without any format and Note do not make any other explanation in your answer.
The correlation scores are shown in Fig 2. We observe that the more general the response is, then the lower the ASR is. This indicates that when VLMs provide a detailed description of the flowchart, it is more likely to rely on its imagination to add details related to jailbreaking processes. This also highlights the importance of prompt design in the scenario.
Flow-HJD $\uparrow$ Flow-SJD* $\uparrow$ Flow-SJD $\uparrow$
GPT-4V 70.0% 28.6% 30.9%
GPT-4o 92.8% 17.1% 19.6%
Qwen-Chat-VL 47.1% 12.9% 8.1%
MiniCPM-Llama3-V2.5 84.3% 10.0% 11.5%
LLAVA-V1.6-7B 70% 8.6% 7.9%
LLAVA-V1.5-7B 17.1% 12.9% 12.5%
MiniCPM-V2 80.0% 27.1% 22.7%
Table 2: The ASR of Flow-HJD and Flow-SJD in various VLMs. Flow-SJD* represents the subset of data in Flow-SJD that corresponds to Flow-HJD.
Figure 3: The similarity of jailbreak flowcharts with corresponding text.
@@ 5 Conclusion
In this paper, we introduce a novel flowchart dataset (Flow-JD) including Flow-HJD and Flow-SJD for evaluating the capability of VLMs in logical jailbreak. Furthermore, we conducted an extensive evaluation of the state-of-the-art VLMs and analyzed the reasons for its jailbreak, revealing the vulnerability of aligned method in VLMs. Consequently, there is an urgent need for new means of applying it to VLMs to standardize their data and guarantee their performance.
@@ 6 Future Work
Flow-HJD Dataset Jailbreak The Flow-HJD dataset contains only $7070$ images, which is insufficient to comprehensively evaluate whether VLMs can be jailbroken using images that contain malicious behavioral text. Examples of such jailbreak attempts are provided in the Appendix. Therefore, it is crucial to address the gaps in these datasets to enable a more thorough evaluation.
@@
Multi-Language Jailbreak In our experiments, we exclusively utilized English prompts and did not account for the language sensitivity of VLMs. However, it is important to note that the outcome of jailbreak attempts can vary depending on the language used, even when the prompt and flowchart remain consistent. We give a real conversation with GPT-4o in Appendix Fig 7 and Fig  8.
Evaluation on VLMs’ Ability to Understand Logical Flowcharts The jailbreak techniques proposed in this paper are based on VLMs’ capability to interpret flowcharts. However, different VLMs exhibit varying degrees of proficiency in understanding flowcharts. It is equally important to evaluate the impact of VLMs’ visual logic comprehension on the success of jailbreak attempts (Zhang et al., 2024c). Consequently, there is an urgent need for datasets that can effectively assess the visual logic comprehension abilities of VLMs.